# Rigidity of Gradient Shrinking Ricci Solitons with Bach-like Flatness and Related Variational Formulas

### v1

@@ 10 Variational Origins of U and V
We now connect the tensors U and V with the variational structure of the Riemannian metric. In this section, we identify Bach-like tensors as the Euler–Lagrange gradients of natural quadratic curvature functionals. Throughout, we work in the standard transverse–traceless (TT) gauge, imposing for a variation $h$
@@
$\mathcal{F}_{\alpha,\beta}(g):=-\,\alpha\int_{M}|W_{g}|^{2}\,dV_{g}\;+\;\frac{\alpha/3-\beta}{2}\int_{M}R_{g}^{2}\,dV_{g}\,.$
@@
Using $\mathrm{B}=\tfrac{1}{2}\,\mathrm{U}+\tfrac{1}{6}\,\mathrm{V}$,
@@
$\delta^{2}F[h,h]=\int_{M}\Big[-\,\mathrm{R}\,|\nabla h|^{2}\;-\;2\mathrm{R}\,\mathrm{R}_{ikjl}\,h_{ij}h_{kl}\;+\;2\mathrm{R}\,\mathrm{R}_{ik}\,h_{ij}h_{kj}\;+\;2\langle\mathrm{Rc},h\rangle^{2}\;+\;\Big(\frac{1}{2}\mathrm{R}^{2}-2\,\Delta\mathrm{R}\Big)|h|^{2}\Big]\,dV\,.\;$
@@
$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=\int_{M}\!\Big\langle\big(\alpha\,\Delta_{L}^{2}+\frac{1}{2}\mathrm{R}\,(\beta-\tfrac{4}{3}\alpha)\,\Delta_{L}+\mathrm{R}^{2}(\tfrac{5}{36}\alpha-\tfrac{1}{4}\beta)\big)h,\;h\Big\rangle dV.$
@@
$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=\alpha\int_{M}\langle\Delta_{L}^{2}h,h\rangle\,dV=\alpha\int_{M}|\Delta h|^{2}\,dV\quad(\text{TT}),$

### v2

@@ 10 Variational Origins of U and V
The classical Bach tensor arises naturally in a number of ways. One approach is to start with the conformal Einstein equations and then study integrability conditions via divergence operations. A necessary condition to be conformally Einstein is for the Bach tensor $\mathrm{B}_{ij}$ to vanish. The approach taken by R. Bach’s [bach:1921] was to consider the first variation of the quadratic Riemannian functional
$\mathrm{S}_{W}=\int|\mathrm{W}|^{2}\,\,\mathrm{d}V_{g}.$
As can be seen in A. Besse [besse:1987], one has
$\mathrm{B}\propto\mathrm{grad}\left(\int|\mathrm{W}|^{2}\,\mathrm{d}V_{g}\right),$
so the Bach tensor is a critical point of a Riemannian functional whose energy depends on how much a metric deviates from being conformally flat. In this section, we identify Bach-like tensors as the Euler–Lagrange gradients of natural quadratic curvature functionals. We will work in the standard transverse–traceless (TT) gauge for the second variation, imposing for a variation $h$
@@
It is worth recalling the Lichnerowicz Laplacian as it will appear in our second variation formulas. This operator $\Delta_{L}$ acts on symmetric 2-tensors via the rule
$(\Delta_{L}\,h)_{ij}=(\nabla^{*}\nabla\,h)_{ij}+2\,\mathrm{R}_{ikjl}\,h_{kl}-\mathrm{R}_{ik}\,h_{kj}-\mathrm{R}_{jk}\,h_{ki}.$
Here $\nabla^{*}$ denotes the divergence operator or formal adjoint of the covariant derivative $\nabla$, and $\nabla^{*}\nabla$ is the rough Laplacian. Note that sign conventions for the Laplacian vary, but we will take the convention that
$\nabla^{*}\nabla\,h=-\nabla^{k}\nabla_{k}\,h.$
@@
$\mathcal{F}_{\alpha,\beta}(g):=-\,\alpha\int_{M}|\mathrm{W}_{g}|^{2}\,\mathrm{d}V_{g}\;+\;\frac{\alpha/3-\beta}{2}\int_{M}\mathrm{R}_{g}^{2}\,\mathrm{d}V_{g}\,.$
@@
These standard variational identities for quadratic functionals are well-established [e.g., [besse:1987]]. Using $\mathrm{B}=\tfrac{1}{2}\,\mathrm{U}+\tfrac{1}{6}\,\mathrm{V}$,
@@
$\delta^{2}F[h,h]=\int_{M}\Big[-\,\mathrm{R}\,|\nabla h|^{2}\;-\;2\mathrm{R}\,\mathrm{R}_{ikjl}\,h_{ij}h_{kl}\;+\;2\mathrm{R}\,\mathrm{R}_{ik}\,h_{ij}h_{kj}\;+\;2\langle\mathrm{Rc},h\rangle^{2}\;+\;\Big(\frac{1}{2}\mathrm{R}^{2}-2\,\Delta\mathrm{R}\Big)|h|^{2}\Big]\,\,\mathrm{d}V\,.\;$
@@
$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=\int_{M}\!\Big\langle\big(\alpha\,\Delta_{L}^{2}+\frac{1}{2}\mathrm{R}\,(\beta-\tfrac{4}{3}\alpha)\,\Delta_{L}+\mathrm{R}^{2}(\tfrac{5}{36}\alpha-\tfrac{1}{4}\beta)\big)h,\;h\Big\rangle\,\mathrm{d}V.$
@@
$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=\alpha\int_{M}\langle\Delta_{L}^{2}h,h\rangle\,\,\mathrm{d}V=\alpha\int_{M}|\Delta h|^{2}\,\,\mathrm{d}V\quad(\text{TT}),$